OpenStudy (anonymous):

A hollow ball with mass 0.5 kg and radius 10 com rolls down a 3 m high ramp without sliding. At the end of the ramp is an incline which launches the ball at 30 degrees. What is the maximum hight that the ball reaches?

OpenStudy (vincent-lyon.fr):

Do you have an idea how you can solve that problem?

OpenStudy (vincent-lyon.fr):

I find 22.5 cm

OpenStudy (anonymous):

the answer is 0.45m but idk why....

OpenStudy (anonymous):

The total mechanical energy of the ball = KE(rotational) + KE(linear). The moment of inertial of a hallow sphere is I = (2/3)mr^2. If KE(rotational) = (1/2)I(omega)^2 where v=(omega)r , you should be able to find the velocity off the ramp. Then from the Kinetic Laws of Motion find the maximum height.

OpenStudy (vincent-lyon.fr):

@a25526598 Can you show us your work? The method described by shawmoes is the right one. First thing we need is the velocity of the ball's centre of mass when it leaves the ramp, with respect to g and ho = 3m